OpenStudy (anonymous):

An object travels half its total path in the last second of its fall from rest. Find the following: A) Total time it falls B) The height of its fall

OpenStudy (anonymous):

Nice problem. Distance s to fall in time t, from rest, is given by s = gt^2. So the problem is: s(t-1) = 1/2 s(t) g(t-1)^2 = 1/2 g t^2 (g = 9.8 m/s^2). Solve for t. Substitute into gt^2 to find s.

OpenStudy (anonymous):

Nice and neat answer from Carl_Pham.